OpenStudy (sheerfire):

Help. The shell of a chicken's egg makes up 5% of the mass of an average egg. An average egg has a mass of 50 g. Assume the egg shell is pure calcium carbonate. How many complete chicken's egg shells would be needed to neutralise 50 cm3 of 2.0 mol dm-3 ethanoic acid?

OpenStudy (jebonna):

You should first write the equation of calcium carbonate + ethanoic acid: 2CH3COOH + CaCO3 -> Ca(CH3CO2)2 + H2O + CO2 Then, work out 5% of 50g to get the mass of the egg shell. You can do this by doing: 50/100 = 0.5g (1%) and then 0.5g x 5 = 2.5g (5%) So the egg shell is 2.5g. So now we should work out the moles of ethanoic acid using this formula: Moles = Concentration (mol dm-3) x Volume (dm3) 2.0 x (50cm3/1000)dm3 = 0.1 mol of ethanoic acid (You have to divide 50cm3 by 1000 to change cm3 into dm3) So, looking at the stoichiometry in the equation we wrote at the beginning, there are TWO lots of ethanoic acid, and ONE lot of calcium carbonate. We need the moles of calcium carbonate (to find the mass in grams), but we have the moles of TWO lots of a substance (number of moles will be the same for each compound in the equation). To find the moles of calcium carbonate, as we have the moles for two lots of something and we need it for one lot of something, we will need to divide 0.1 by 2 to get the moles for one lot of a substance in our equation. 0.1 / 2 = 0.05mol of calcium carbonate Now we can find the mass in grams of calcium carbonate so we know how many eggs we need to match the number of grams needed to neutralise ethanoic acid. To find the mass of a substance you will need to use this equation: Mass (g) = Moles (mol) x Molar mass (g mol-1) First work out the molar mass of calcium carbonate: Ca = 40.1 C = 12 O = 16 40.1 + 12 + (16 x 3) = 100.1g mol-1 Now you can work out the mass in grams: 0.05mol x 100.1g mol-1 = 5g of calcium carbonate needed (Rounded to a whole number) So now you know you need 5g of calcium carbonate to neutralise 50cm3 ethanoic acid. To find out how many eggs you need, you'll need to divide the mass of calcium carbonate needed, by the mass of egg shell to find how many eggs you'll need in total. 5g / 2.5g = 2 eggs So, you will need 2 egg shells (eggs) to totally neutralise ethanoic acid. I hope this helped! :)

OpenStudy (sheerfire):

Tysm☺

OpenStudy (jebonna):

No problem :)